OpenStudy (anonymous):

Describe the graph of the solution in words. |3x - 2| < 7

OpenStudy (anonymous):

3x - 2 < 7 3x < 7 + 2 3x < 9 x < 3

OpenStudy (anonymous):

What would the graph look like

OpenStudy (anonymous):

Distribute the 3 and the 2 3(2) - 3(3x) = -7 - 2(x) - 2(-3) yields 6 - 9x = -7 - 2x + 6 ... Combine like terms 2x - 9x = -7 + 6 - 6 -7x = -7 ... Divide by -7

OpenStudy (anonymous):

so what would the graph look like

OpenStudy (anonymous):

??